The product of twice a number increased by 5 and the number squared is 18.
This has to be written out in an equation and then solved. So far I have...
(2n+5)[(n)(n)]=18
then I can get it as simplified as
2(n)(n)(n)+5(n)(n)=18
is this right? (sorry I can't get my computer to type powers)
PLEASE HELP ME!

Answer by checkley71(8403)

(2N+5)(N~2)=18 OR 2N~3+5N~2=18 OR N=1.5 (TRIAL & ERROR METHOD)
PROOF 2(1.5)~3+5(1.5)~2=18 OR 2(3.375)+5(2.25)=18 OR 6.75+11.25=18 OR 18=18